OpenStudy (anonymous):

Solve 4x + 7 = 6x – 1

OpenStudy (hba):

put the terms including x on one side and the other on one side and show me

OpenStudy (anonymous):

Like combined like terms??

OpenStudy (campbell_st):

subtract 4x from both sides 7 = 2x - 1 add 1 to both sides 8 = 2x solve for x

OpenStudy (anonymous):

x=0.25

OpenStudy (campbell_st):

not quite what if I wrote it 2x = 8

OpenStudy (anonymous):

ummm.... sorry i wrote the problem wrong it is 4^(x+7)=6^(x-1)

OpenStudy (campbell_st):

lol...

OpenStudy (hba):

use log

OpenStudy (anonymous):

Do you know how to solve that???

OpenStudy (hba):

yes

OpenStudy (hba):